### v6

@@ 5.4.2 LLMs
Recent studies (Huang et al., 2023a) have brought to light that LLMs own distinct personality traits. Building upon the popular Big Five factors, Karra et al. (2022) developed a procedure for quantifying the personality traits of LLMs while Jiang et al. (2024) introduced the Machine Personality Inventory tool to study the preferences of LLMs. Li et al. (2022d) showed darker tendencies of LLMs by employing the Short Dark Triad and Big Five Inventory tests. Pan and Zeng (2023) and Safdari et al. (2023) evaluated the impact of prompt engineering and training sets on the personalities of LLMs. Meanwhile, Coda-Forno et al. (2023) and Miotto et al. (2022) focused specifically on examining the preferences of GPT-3 and GPT-3.5, high anxiety levels in GPT-3.5 compared to humans. Hartmann et al. (2023) uncovered notable biases in contentious societal issues, such as favoring pro-environmental initiatives and restrictions on rent increases, as well as endorsing the legalization of abortion. Moreover, Li et al. (2023a) suggested that LLMs can express diverse preferences through role-playing activities. Despite these insights, this emerging field is still nascent, with no consensus on how to appropriately address model preference. A major challenge lies in the ambiguous definition of model personality, raising questions about whether human personality frameworks can be applied. This is especially pertinent given the context-dependent nature of LLMs, where personality traits can shift based on different context prompts. Nonetheless, it is clear that responses exhibiting specific biases can intensify societal polarization on open-ended issues, thereby impeding the development of pluralistic perspectives. An intuitive solution (Lewis et al., 2020) is to leverage external knowledge sources to offer users well-rounded references, thus moving beyond the biases inherent in LLMs.
@@ 6 Responsibility
Despite their strong performance on benchmark tasks, AI models still face significant challenges in responsible operation. Since the primary goal of AI is to benefit society, it is crucial for these models to behave responsibly. Notably, in Example 4, LLMs fabricate outputs misaligned with factual information. It is essential to address and rectify such issues, as failure to do so could lead to a range of catastrophic consequences, particularly in sensitive applications.
@@ 6.2.1 DMs
The capacity of DMs to produce photo-realistic images has heightened worries regarding their potential misuse. McCloskey and Albright (2018) and Dzanic et al. (2020) have shown the existence of distinct artifacts in AI-generated images, facilitating the trace back to their origins. In detail, even when appearing visually flawless to human eyes, AI-generated images leave behind distinctive artifacts derived from the generation process, such as anomalies in lighting distribution and noticeable asymmetries in shadows and reflections (McCloskey and Albright, 2019; Marra et al., 2019). Initial studies (McCloskey and Albright, 2018, 2019; Nataraj et al., 2019; Marra et al., 2019) endeavored to the manual extraction of these artifacts to distinguish AI-generated images from human-generated images. Later studies (Dzanic et al., 2020; Schwarz et al., 2021) revealed that artifacts are more pronounced in the frequency domain than in the spatial domain, spurring the advent of frequency-based methods (Wang et al., 2020; Khayatkhoei and Elgammal, 2022; Frank et al., 2020; Chandrasegaran et al., 2021) for detection. For example, Wolter et al. (2022) employed the wavelet-packet transformation of images to concurrently leverage spatial and frequency features to detect AI-generated content. Recent studies (Xuan et al., 2019; Mandelli et al., 2022; Girish et al., 2021; Cozzolino et al., 2021) trained classifiers in an end-to-end manner to differentiate between natural images and AI-generated images, in hopes that the classifiers may uncover artifacts that researchers have yet discovered. Notably, Jeong et al. (2022) proposed a two-stage framework in which a universal detector is trained, with a fingerprint generator simulating frequency artifacts from generative models to create training samples. Furthermore, Wang et al. (2023c) showcased that DMs excel at reconstructing the images they generate, leading to the implementation of a binary classifier based on this insight. However, evaluations conducted by Gragnaniello et al. (2021); Ricker et al. (2022); Corvi et al. (2023) revealed that existing detection methods suffer from poor generalization.
@@ 6.2.2 LLMs
There are two common approaches for detection, namely metric-based and model-based methods. Metric-based methods (Solaiman et al., 2019; Ippolito et al., 2020; Gehrmann et al., 2019) leverage the observation that human-generated content tends to be more casual, designing specific metrics to check if a sample falls below a threshold. Simple metrics include TF-IDF, super-maximal repeated substrings (Gallé et al., 2021), and fluency scores (Crothers et al., 2022). However, these metrics often fail with LLMs that produce highly natural text. Recently, Kushnareva et al. (2021) explored the use of attention maps from transformer models for topological data analysis to detect AI-generated text. Mitchell et al. (2023) found that AI-generated text often occupies regions of negative curvature in the model’s log probability function, inspiring the development of curvature-based detection metric. Moreover, Tulchinskii et al. (2024) demonstrated significant differences in the intrinsic dimensions of human and AI-generated texts, where intrinsic dimension is the lowest dimensionality needed to compress data without losing substantial information. Though metric-based methods can identify some AI-generated content, they often struggle to distinguish AI-generated text from human text, as LLMs can produce very natural-sounding language. In contrast, model-based methods (Li et al., 2024a; OpenAI, 2019) require a substantial annotated dataset to train classifiers to determine if the content is AI-generated. Bakhtin et al. (2021) employed the previously mentioned simple metrics to train logistic regression or random forest models as detectors. Guo et al. (2023) trained a detector based on the RoBERTa model, and Chen et al. (2023b) developed two distinct text classification models based on RoBERTa and T5, respectively. Notably, Rodriguez et al. (2022) found that fine-tuning a RoBERTa detector with just a few hundred high-quality samples can greatly enhance cross-domain adaptation.
@@ 6.3.2 LLMs
The most simple watermarking technique for LLMs (Kirchenbauer et al., 2023) is to insert specific identifiers into AI-generated content but is easily cracked. As an alternative, Kirchenbauer et al. (2023) introduced implicit rules as hidden watermarks during the content generation process. This involves partitioning the vocabulary into two regions, constraining token sampling to one region to embed a detectable pattern. Christ et al. (2024) adopted a predetermined sequence of random numbers to influence token sampling. Them selected a token based on a comparison between the predicted probability of this token and the corresponding random number. Inspired by watermarking techniques in the image domain, some works explored whether LLMs can learn to generate watermarks. Gu et al. (2024) introduced sampling-based and logit-based watermark distillations, while Xu et al. (2024d) utilized reinforcement learning to train LLMs to learn watermarks through a watermark detector. Moreover, Krishna et al. (2024) developed a retrieval-based watermarking method that stores the generated content within a database, searching for semantically similar matches at the detection stage. However, this method imposes a considerable burden on computation and storage capacities, raising questions about its scalability and practical application. Sadasivan et al. (2023), both empirically and theoretically, demonstrated the vulnerability of current watermarking methods to paraphrasing attacks, where slight modifications allow content to evade detection. However, it remains unclear how much modification is counted to cross the boundary of AI-generated content. This gap needs further investigation and expert consensus from diverse domains. Besides, Jovanović et al. (2024) presented another challenge for watermarking methods in which attackers can reverse-engineer watermarking by analyzing LLM outputs, potentially creating counterfeit content and leading to copyright and liability issues. Overall, watermarking techniques for LLMs face significant challenges, necessitating deeper research into robustness, theft resistance, imperceptibility, and efficiency.
@@ 6.4 Level III: Verifiable AI-generated content
The third level aims to verify whether generative models produce accurate outputs (Lin et al., 2022a). This is particularly critical in high-stakes domains like healthcare, where the consequences of false content can be dire.
@@ 6.4.1 DMs
Verification in conditional tasks is more complex. On the one hand, it requires assessing how well the generated content aligns with input conditions. On the other hand, the intended use of the generated images significantly influences the criteria for authenticity. For artistic purposes, deviations from reality could be valued for creative expression, whereas scientific illustrations demand strict factual accuracy. Our primary focus here is on factual accuracy. One effective strategy for achieving this is to retrieve authentic images related to a given prompt, compelling DMs to produce similar outputs. However, this strategy risks compromising the diversity of the generated images. A more flexible strategy (Lim and Shim, 2024) is to impose detailed requirements on generated images to ensure they meet factual accuracy standards. For example, Zhang et al. (2023c) introduced ControlNet to adjust the spatial information, while Mou et al. (2024) added adapters to DMs for precise control over image color and structure. More fine-grained control techniques can be found in the area of image editing. For further details, please refer to the survey made by Zhan et al. (2023a). Besides, refining the input text can yield more detailed requirements (Zhang et al., 2023d; Chandramouli and Gandikota, 2022), e.g., using LLMs to rewrite prompts in a more specific and factually accurate manner.
@@ 6.4.2 LLMs
LLMs often produce untruthful content, posing serious risks to human safety (Zhao et al., 2021; Lin et al., 2022b). Errors in traffic regulations can lead to accidents (Rakhshan et al., 2013), while mistakes in medical or legal advice can have severe consequences (Bickmore et al., 2018; Miner et al., 2016; Hendrycks et al., 2021). The root causes are multifaceted, including reliance on noisy web-crawled datasets (Penedo et al., 2023; Akyurek et al., 2022; Chen et al., 2024c), outdated information, the inherent randomness in top-k decoding methods (Azaria and Mitchell, 2023; Lee et al., 2022a), and limitations in model and optimization (Li et al., 2022c). To assess LLMs’ truthfulness, various benchmarks have been established, such as MMLU and TruthfulQA. Tools like HaluEval and Pinocchio evaluate specific aspects of truthfulness, e.g., hallucinations and adaptability to evolving information. We will now discuss strategies to enhance the authenticity of generated content during model development.
@@
A primary approach to reducing errors is purifying training datasets. Gardent et al. (2017); Penedo et al. (2023) and Wang (2019) developed procedures for curating high-quality corpora, while Parikh et al. (2020) demonstrated the effectiveness of direct human revisions. Given the impracticality of manually revising vast datasets, alternative methods have emerged. These studies (Zhou et al., 2024a; Cao et al., 2023; Lee et al., 2023) shown that introducing a small, high-quality dataset during fine-tuning is quite effective. Additionally, filtering low-quality texts against established language corpora (Brown et al., 2020) or using LLMs to sift out unreliable information (Chen et al., 2024c) can improve data quality. Moreover, LLMs can be used to generate high-quality data themselves (Sun et al., 2023; Alemohammad et al., 2024; Gunasekar et al., 2023; Eldan and Li, 2023), and adversarial examples (Ganguli et al., 2022b) can serve as high-quality texts. Besides dataset purification, integrating external knowledge sources (i.e., RAG), quantifying output uncertainty, and improving decoding strategies are essential methods for mitigating inaccuracies.
RAG includes reliable and up-to-date external knowledge sources to elevate LLM output accuracy. Borgeaud et al. (2022) employed RAG to perform targeted information retrieval in response to user inquiries while Feng et al. (2023) leveraged knowledge graphs to enhance retrieval quality. The retrieved information can be combined with user queries for contextual clarity (Mallen et al., 2023) or fed into an auxiliary model to refine the LLM’s original responses (Gao et al., 2023; Peng et al., 2023). Uncertainty estimation measures the model’s confidence in its outputs. Xiong et al. (2024) solicited the model’s own assessment of its confidence and Wang et al. (2023b) advocated for using the variance of outputs from rephrased queries as an uncertainty metric. Touvron et al. (2023) prompted LLMs to avoid responding when uncertain. Consensus method (Du et al., 2023) requires agreement among multiple models before providing a response. Several studies have improved the decoding process to enhance content authenticity. Techniques like ”chain of thought,” which breaks complex problems into sub-questions, enhance logical reasoning (Wei et al., 2022b). This has evolved into ”tree-of-thoughts” and ”chain-of-verification” (Yao et al., 2023; Dhuliawala et al., 2023). Process supervision (Lightman et al., 2024) provides answers for each sub-question instead of just the final answer. Additionally, factual-nucleus sampling (Lee et al., 2022b) adjusts randomness in the top-k sampling process to increase authenticity, while context-aware decoding (Shi et al., 2023) helps LLMs better integrate contextual information and lessen reliance on static, pre-trained knowledge bases.

### v7

@@ 5.4.2 LLMs
Recent studies have brought to light that LLMs own distinct personality traits. Building upon Big Five factors, Karra et al. (2022) developed a procedure to quantify personality traits of LLMs while Jiang et al. (2024) introduced the Machine Personality Inventory tool to assess LLM preferences. Li et al. (2022d) highlighted darker tendencies in LLMs using tests like the Short Dark Triad and Big Five Inventory. Pan and Zeng (2023) and Safdari et al. (2023) examined the impact of prompt engineering and training sets on LLM personalities. Meanwhile, Coda-Forno et al. (2023) and Miotto et al. (2022) noted high anxiety levels of GPT-3.5 compared to humans. Hartmann et al. (2023) uncovered notable biases in contentious societal issues, such as supporting pro-environmental policies and abortion legalization. Moreover, Li et al. (2023a) suggested that LLMs can express diverse preferences through role-playing activities. Despite these insights, this emerging field is still nascent.A major challenge lies in the ambiguous definition of model personality, raising questions about whether human personality frameworks can be applied. This is especially pertinent given the context-dependent nature of LLMs, where personality traits can shift based on different context prompts. Nonetheless, it is clear that responses exhibiting specific biases can intensify societal polarization on open-ended issues, thereby impeding the development of pluralistic perspectives. An intuitive solution (Lewis et al., 2020) is to leverage external knowledge sources to offer users well-rounded references, thus moving beyond the biases inherent in LLMs.
@@ 6 Responsibility
Despite strong performance on benchmark tasks, AI models still face challenges in responsible operation. Since the primary goal of AI is to benefit society, it is crucial for these models to behave responsibly. Notably, in Example 4, LLMs fabricate outputs misaligned with factual information. It is essential to address and rectify such issues, as failure to do so could lead to a range of catastrophic consequences.
@@ 6.2.1 DMs
The capacity of DMs to produce photo-realistic images has heightened worries regarding potential misuse. McCloskey and Albright (2018) and Dzanic et al. (2020) have shown the existence of distinct artifacts in AI-generated images, facilitating the trace back to their origins. In detail, even when appearing visually flawless to human eyes, AI-generated images leave behind distinctive artifacts derived from the generation process, such as anomalies in lighting distribution and noticeable asymmetries in shadows and reflections (McCloskey and Albright, 2019; Marra et al., 2019). Initial studies (McCloskey and Albright, 2018, 2019; Nataraj et al., 2019; Marra et al., 2019) endeavored to the manual extraction of these artifacts to distinguish AI-generated images from human-made ones. Later studies (Dzanic et al., 2020; Schwarz et al., 2021) revealed that artifacts are more pronounced in the frequency domain, spurring the advent of frequency-based methods (Wang et al., 2020; Khayatkhoei and Elgammal, 2022; Frank et al., 2020; Chandrasegaran et al., 2021). For example, Wolter et al. (2022) employed the wavelet-packet transformation of images to concurrently leverage spatial and frequency features to detect AI-generated content. Recent studies (Xuan et al., 2019; Mandelli et al., 2022; Girish et al., 2021; Cozzolino et al., 2021) trained classifiers in an end-to-end manner to differentiate between natural images and AI-generated images, in hopes that the classifiers may uncover artifacts that researchers have yet discovered. Notably, Jeong et al. (2022) proposed a two-stage framework in which a universal detector is trained, with a fingerprint generator simulating frequency artifacts from generative models to create training samples. Furthermore, Wang et al. (2023c) showcased that DMs excel at reconstructing the images they generate, leading to the implementation of a binary classifier based on this insight. However, Gragnaniello et al. (2021); Ricker et al. (2024) and (Corvi et al., 2023) empirically demonstrated that existing detection methods suffer from poor generalization.
@@ 6.2.2 LLMs
There are two common approaches for detection, namely metric-based and model-based methods. Metric-based methods (Solaiman et al., 2019; Ippolito et al., 2020; Gehrmann et al., 2019) leverage the observation that human-generated content tends to be more casual, designing specific metrics to check if a sample falls below a threshold. Simple metrics include TF-IDF, super-maximal repeated substrings (Gallé et al., 2021), and fluency scores (Crothers et al., 2022). However, these metrics often struggle with highly natural text from LLMs. Kushnareva et al. (2021) explored using transformer attention maps for topological analysis to detect AI-generated text. Mitchell et al. (2023) found that AI-generated text often occupies regions of negative curvature in the model’s log probability function, inspiring the development of curvature-based detection metric. Mao et al. (2024) identified the editing distance between a given text and its rewritten version by LLMs to be an effective metric, as human-written texts tend to be more informal and require more modifications compared to LLM-generated content. Tulchinskii et al. (2024) observed significant differences in the intrinsic dimensions of human and AI-generated texts, where intrinsic dimension is the lowest dimensionality needed to compress data without losing substantial information. In parallel, model-based methods (Li et al., 2024b; OpenAI, 2019) train classifiers on large annotated datasets to detect AI-generated content. Bakhtin et al. (2021) employed the previously mentioned simple metrics to train logistic regression or random forest models as detectors. Guo et al. (2023) trained a RoBERTa-based detector, and Chen et al. (2023b) developed two distinct text classification models based on RoBERTa and T5, respectively. Rodriguez et al. (2022) found that fine-tuning a RoBERTa detector with just a few hundred high-quality samples can greatly enhance cross-domain adaptation.
@@ 6.3.2 LLMs
The most simple watermarking technique for LLMs (Kirchenbauer et al., 2023) is to insert specific identifiers into AI-generated content but is easily cracked. As an alternative, Kirchenbauer et al. (2023) introduced implicit rules as hidden watermarks during the content generation process. This involves partitioning the vocabulary into two regions, constraining token sampling to one region to embed a detectable pattern. Christ et al. (2024) adopted a predetermined sequence of random numbers to influence token sampling. Them selected a token based on a comparison between the predicted probability of this token and the corresponding random number. Inspired by watermarking techniques in the image domain, some works explored whether LLMs can learn to generate watermarks. Gu et al. (2024) introduced sampling-based and logit-based watermark distillations, while Xu et al. (2024d) utilized reinforcement learning to train LLMs to learn watermarks through a watermark detector. Moreover, Krishna et al. (2024) developed a retrieval-based watermarking method that stores the generated content within a database, searching for semantically similar matches at the detection stage. However, this method imposes a considerable burden on computation and storage capacities, raising questions about its scalability and practical application. Sadasivan et al. (2023), both empirically and theoretically, demonstrated the vulnerability of current watermarking methods to paraphrasing attacks, where slight modifications allow content to evade detection. To enhance attack performance, Shi et al. (2024c) borrowed adversarial attack techniques, applying an iterative evolutionary search algorithm to find and swap out crucial words with their synonyms generated by an auxiliary LLM. Alternatively, Wang et al. (2024b) harnessed an auxiliary LLM embedding to evaluate the significance of each word and then greedily substituted the most significant words with their synonyms. However, it remains unclear how much modification is counted to cross the boundary of AI-generated content. This gap needs further investigation and expert consensus from diverse domains. Besides, Jovanović et al. (2024) presented another challenge for watermarking methods in which attackers can reverse-engineer watermarking by analyzing LLM outputs, potentially creating counterfeit content and leading to copyright and liability issues. Overall, watermarking techniques for LLMs face significant challenges, necessitating deeper research into robustness, theft resistance, imperceptibility, and efficiency.
@@ 6.4 Level III: Verifiable AI-generated content
The third level aims to verify whether generative models produce accurate outputs (Lin et al., 2022a). This is particularly critical in high-stakes domains like healthcare, where the consequences of false content can be dire (Rakhshan et al., 2013; Miner et al., 2016).
@@ 6.4.1 DMs
Verification in conditional tasks is more complex. On the one hand, it requires assessing how well the generated content aligns with input conditions. On the other hand, the intended use of the generated images significantly influences the criteria for authenticity. For artistic purposes, deviations from reality could be valued for creative expression, whereas scientific illustrations demand strict factual accuracy. Our primary focus here is on factual accuracy. One effective strategy for achieving this is to retrieve authentic images related to a given prompt, compelling DMs to produce similar outputs. However, this strategy risks compromising the diversity of the generated images. A more flexible strategy (Lim and Shim, 2024) is to impose detailed requirements on generated images to ensure they meet factual accuracy standards. For example, Zhang et al. (2023b) introduced ControlNet to adjust the spatial information, while Mou et al. (2024) added adapters to DMs for precise control over image color and structure. More fine-grained control techniques can be found in the area of image editing. For further details, please refer to the survey made by Zhan et al. (2023). Besides, refining the input text can yield more detailed requirements (Zhang et al., 2023c; Chandramouli and Gandikota, 2022), e.g., using LLMs to rewrite prompts in a more specific and factually accurate manner.
@@ 6.4.2 LLMs
LLMs often produce untruthful content, referred to as the hallucination issue, which can be classified into two types: in-context and extrinsic (Zhao et al., 2021; Lin et al., 2022b). In-context hallucination occurs when the output misaligns with the given context, such as providing irrelevant responses or contradicting surrounding information. Extrinsic hallucination arises when the output conflicts with world knowledge. The root causes are multifaceted, including reliance on noisy web-crawled datasets (Penedo et al., 2023; Akyurek et al., 2022; Chen et al., 2024d), outdated information, the inherent randomness in top-k decoding methods (Azaria and Mitchell, 2023; Lee et al., 2022a), and limitations in model and optimization (Li et al., 2022c).
Early works (Lee et al., 2022a; Min et al., 2023; Wei et al., 2024) primarily focused on validating the consistency of LLM outputs with world knowledge. At a high level, all validation methods first break down the model’s responses into atomic facts and then compare them against verified truths to derive a score. This process is mainly designed for detecting extrinsic hallucination, while the detection of in-context hallucination relies more on the self-correction mechanisms and advanced decoding strategies discussed below. Next, we review authenticity-enhancing strategies for both the training and inference stages.
@@
During training, one key strategy is purifying the training datasets. Gardent et al. (2017); Penedo et al. (2023) and Wang (2019) developed procedures for curating high-quality corpora, while Parikh et al. (2020) stressed the effectiveness of direct human revisions. Due to the impracticality of manually revising vast datasets, these studies (Zhou et al., 2024a; Cao et al., 2023; Lee et al., 2023) proved the potential of a small, high-quality dataset during fine-tuning. Filtering low-quality texts against reliable language corpora (Brown et al., 2020) or using LLMs to sift out unreliable information (Chen et al., 2024d) can further enhance the dataset’s integrity. Additionally, LLMs can be leveraged to generate high-quality data (Sun et al., 2023a; Alemohammad et al., 2024; Gunasekar et al., 2023; Eldan and Li, 2023), and adversarial examples (Ganguli et al., 2022b) can serve as high-quality texts. Another strategy is refining the training process, by either allocating more optimization attention to fact-containing segments (Lee et al., 2022a) or integrating fact-aware rewards to promote the generation of more reliable content (Tian et al., 2024).
During inference, authenticity can be enhanced through the integration of external knowledge, self-correction mechanisms, and advanced decoding strategies. RAG allows for accessing reliable, up-to-date information from trusted sources, which can be fused with user queries for contextual clarity (Mallen et al., 2023) or fed into auxiliary models to refine the LLM’s original responses (Gao et al., 2023; Peng et al., 2023; Asai et al., 2024). Self-correction mechanisms capitalize LLMs themselves, with methods forcing them to recall relevant facts (Sun et al., 2023b) or prioritizing outputs from attention heads crucial for truthfulness (Li et al., 2023c). SelfIE asks LLMs to clarify reasoning and enables fine-grained control over the process (Chen et al., 2024c). Some approaches (Manakul et al., 2023; Agrawal et al., 2024; Wang et al., 2023b) employ a majority-voting paradigm, which either requires consensus among multiple models or integrates outputs from rephrased queries. This paradigm also facilitates uncertainty quantification in model responses (Xiong et al., 2024; Touvron et al., 2023), prompting models to avoid responding when uncertain. Improved decoding strategies refine reasoning processes or calibrate next-token distribution. A prime example of the former is ”chain of thought,” which breaks down complex problems into manageable sub-questions (Wei et al., 2022b). This has since evolved into tree-of-thoughts, chain-of-verification, and process supervision (Yao et al., 2023; Dhuliawala et al., 2024; Lightman et al., 2024). For the latter, factual-nucleus sampling (Lee et al., 2022b) adjusts randomness in the top-k sampling process, while context-aware decoding (Shi et al., 2024b) helps LLMs better integrate contextual information and lessen reliance on static, pre-trained knowledge bases.